# PoI: A Filter to Extract Pixel of Interest from Novel View Synthesis for Scene Coordinate Regression

### v3

@@ Abstract
Novel View synthesis (NVS) techniques, notably Neural Radiance Fields (NeRF) and 3D Gaussian Splatting (3DGS), can augment camera pose estimation by extending and diversifying training data. However, the images generated by these methods are often plagued by spatial artifacts such as blurring and ghosting, undermining their reliability as training data for camera pose estimation. This limitation is particularly critical for Scene Coordinate Regression (SCR) methods, which aim at pixel-level 3D coordinate estimation, because rendering artifacts directly lead to estimation inaccuracies. To address this challenge, we propose a dual-criteria filtering mechanism that dynamically identifies and discards suboptimal pixels during training. The dual-criteria filter evaluates two concurrent metrics: (1) real-time SCR reprojection error, and (2) gradient threshold, across the coordinate regression domain. In addition, for visual localization problems in sparse input scenarios, it will be even more necessary to use data generated by NVS to assist the localization task. We design a coarse-to-fine PoI variant using sparse input NVS to solve this problem. Experiments across indoor and outdoor benchmarks confirm our method’s efficacy. It achieves state-of-the-art localization accuracy while maintaining computational efficiency.
Index Terms:
Visual Localization, Novel View synthesis, Scene Coordinate Regression.
@@ I Introduction
Figure 1: Left: Comparison of query and rendered images of the dataset 7Scenes and Cambridge Landmarks, revealing uneven quality within them, with query frames clear and rendered frames blurry or ghosted; Right: Translation error versus training time, where ‘CoodiNet+’ denotes using rendered images as query images for CPR method CoodiNet (LENS in this case); ‘DSAC*+’ and ‘ACE+’ denote the method combines NVS-rendered images and query images as training data for SCR method DSAC* and ACE. ‘ACE+PoI’ denotes our proposed PoI method (ACE-based); Analysis reveals that directly adding rendered data to the training set will increase training time to some extent, but performance will decrease for the SCR method. On the other hand, our PoI approach can improve the performance with an acceptable time increase.
@@
Although artifacts exist in rendered images as shown in the left part of Figure 1, the CPR-based NRP methods achieve effectiveness mainly because they involve N-to-1 prediction (1 prediction for N pixels) and are more affected by the overall quality of the images than particular pixels. In contrast, SCR-based methods rely on accurate N-to-N prediction (N predictions for N pixels) and are particularly susceptible to the quality of rendered pixels. As shown in the right part of Figure 1, after extending the training dataset with synthetic data through NVS, the CPR method shows significant improvement, while the performance of SCR declines, accompanied by a notable increase in training time. Directly training the SCR model with raw rendered images proves less effective than CPR methods. As a result, very few studies can effectively combine NVS with the SCR framework to expand the training dataset.
@@ II Related Work
Camera Pose Regression The CPR methods, i.e., to regress the camera pose from the given image directly, are the most naive ideas in learning-based methods [20, 21, 22, 23, 24, 25, 26, 27, 12, 28]. The most straightforward method implicitly uses CNN layers or MLP to represent the image-to-pose correspondence. PoseNet [20] first proposes CPR using pre-trained GoogLeNet as the feature extractor. Then, several works focus on improving CPR through additional modules. Geomapnet [22] estimates the absolute camera poses and the relative poses between adjacent frames. AtLoc [25] uses a self-attention module to extract salient features from the image. Vlocnet++ [24] adds a semantic module to solve the dynamic scene and improve the robustness for blockings and blurs. Marepo [29] first regresses the scene-specific geometry from the input images and then estimates the camera pose using a scene-agnostic transformer. The CPR method has achieved excellent efficiency and framework simplification, but there is still room for improvement in accuracy.
Scene Coordinate Regression The SCR-based method [16, 14, 30, 15, 31, 32, 13, 33] aims to estimate the camera pose in two steps: first, estimate the coordinates of the points in the 3D scene corresponding to the 2D pixels, and then estimate the camera pose based on the 2D-3D correspondence relationship. SCR was initially proposed using the random forest for RGB-D images [16]. Recently, estimating scene coordinates through RGB input has been widely studied. ForestNet[31] compares the benefits of Random Forest (RF) and Neural Networks in evaluating the scene coordinate and camera poses. ForestNet also proposes a novel method to initiate the neural network from an RF. DSAC [14], DSAC++ [30] devise a differentiable RANSAC, and thus the SCR method can be trained in an end-to-end manner. ESAC [15] uses a mixture of expert models (i.e., a gating network) to decide which domain the query belongs to, and then the complex SCR task can be split into simpler ones. DSAC* [13] extends the previous works to applications using RGB-D or RGB images, with/without the 3D models. This means that in the minimal case, only RGB images will be used as the input to DSAC*, just like most CPR methods. More information about the 3D structure will be utilized for most SCR methods than for CPR ones. However, approaches like DSAC* can achieve more accurate estimations even if the input is the same as the CPR method. ACE [18] and GLACE [34] abandon the time-consuming end-to-end supervision module and shuffle all pixels of the scene to improve training efficiency. ACE and GLACE use only RGBs without extra 3D geometry information and achieve comparable accuracy compared with former methods.
@@
Figure 2: Pipeline of our proposed methods: (a) Data Prepare: We first sample a group of synthesized camera poses $P_{novel}$ according to the query training pose $P_{query}$ using Grid Sample. Then, we render the synthesized views $I_{novel}$ based on the sampled poses $P_{novel}$ through the novel view synthesis model. (b) Architecture of PoI module: First, a pre-trained scene-irrelevant backbone is applied to extract the features of the input query photos $I_{query}$ and the synthesized novel images $I_{novel}$. Then, the filter is applied to the features of the rendered images and extracts the features of interest. After that, we combine the query features with the retained features of the novel views and shuffle the pixel-aligned features to get the aggregation. Finally, we estimate the scene coordinates of the pixels using a scene-specific Head. The filtering algorithm is designed based on the reprojection error and the gradient of the estimated scene coordinates.
@@ III Method
In general, the pipeline of our proposed method is shown in Figure 2. For the input query images $I_{query}$ and corresponding camera poses $P_{query}$, we first sample the novel camera pose $P_{novel}$ using Grid Sampling. Then we render novel views $I_{novel}$ through 3DGS. Finally, we use the PoI module to estimate the scene coordinates according to the input $I_{query},I_{novel}$. During the evaluation stage, we use PNP-based RANSAC to infer the camera poses from the scene coordinates.
The following part of this section is arranged as follows: We first introduce the building blocks of this paper, including NVS models, the synthetic pose sampling approach, and the baseline of the SCR regressor in Section III-A; Then we explain the details of grid sample method in Section III-B; We elaborate on the architecture of the PoI method and introduce the details of the filtering strategy used in the PoI method in Section III-C and Section III-D; We finally explain the variant of the PoI method in extreme scenarios of sparse input in Section III-E.
@@ III-A Building Blocks
Backbone: To expedite convergence, we use the pseudo-depth supervision mechanism from ACE for suboptimal pixels exhibiting elevated reprojection errors. Specifically, when a pixel’s reprojection error surpasses a predefined threshold, we initially enforce their convergence toward designated pseudo-depth targets through constrained optimization. After a certain epoch, when the estimated values of these pixels fall within the valid threshold, we use conventional supervision for further training.
III-B Grid Sample Method
Before PoI training, we would sample novel camera poses $P_{novel}$ and render the corresponding images $I_{novel}$ based on the query images $I_{query}$ and the corresponding camera poses $P_{query}$. In existing novel view synthesis supported visual localization, we usually have to balance the novel poses’ diversity and the images’ overall rendering quality. However, since we design the strategy to choose POIs from rendered images, we can prioritize pose diversity over rendering quality during the sampling phase. Therefore, we use a unified sampling method for camera pose translation: grid sampling. The boundaries of the grid are calculated based on the camera pose of the training data; that is, the maximum and minimum values of the grid (x, y, z) are determined by the maximum and minimum values of the translations of all camera poses. We add a random perturbation to the rotation. The original rotation of each grid starts from the closest camera pose to that grid in the training data.
III-C Architecture of PoI
@@ III-D Filtering Strategy in PoI
=$\begin{split}\mathcal{L}&=\begin{cases}\mathcal{L}_{rep}^{query}(i),&\text{if }i\in T\\ \tilde{\omega}\times\mathcal{L}_{rep}^{poi}(i),&\text{if }i\in PoI\end{cases}\\ \tilde{\omega}&=\omega_{max}-\frac{I_{iter}}{N_{iter}}(\omega_{max}-\omega_{min})\end{split}$ (2)
@@
III-E Coarse-to-Fine Strategy for Extreme Sparse Input
@@ IV Experiment
PoI(ours) 0.5/0.15 0.7/0.31 0.6/0.30 0.9/0.26 1.0/0.20 0.7/0.22 2.4/0.70 1.0/0.31
GLPoI(ours) 0.6/0.19 0.8/0.33 0.5/0.33 0.9/0.29 0.8/0.22 0.6/0.19 2.8/0.77 1.0/0.33
@@ IV-A Experiments Setup
PoI (ours) No 25min 16M 18/0.3 23/0.5 5/0.2 9/0.3 27/0.1 16/0.3
GLPoI (ours) No 25min 13M 19/0.3 16/0.4 4/0.2 8/0.3 18/0.1 13/0.3
@@ IV-D Coarse-to-Fine Experiments with Extremely Sparse Input
To further evaluate the effectiveness of our method, we conduct an additional experiment with sparse input as mentioned in Section III-E. The comparison of translation error between PoI and C2F for scene ‘fire’ is shown in Figure 7. We can see that C2F has an advantage when the input data is sparse.
@@
The numerical results are shown in Table IV, case ‘base’ denotes the sparse input circumstances with the baseline model. Case ‘coarse’ is the self-pruning step of our method only using rendered data; we still use grid sampling as the novel pose sampling method. Case ‘c2f’ denotes the fine-tuned results of our proposed method.
@@
$\downarrow$ $\downarrow$ $U_{5\text{cm},5\degree}\uparrow$ $\downarrow$ $\downarrow$ $U_{10cm,5\degree}\uparrow$
@@
$\downarrow$ $\downarrow$ $U_{5cm,5\degree}\uparrow$ $\downarrow$ $\downarrow$ $U_{10cm,5\degree}\uparrow$

### v4

@@ Abstract
Novel View synthesis (NVS) techniques, notably Neural Radiance Fields (NeRF) and 3D Gaussian Splatting (3DGS), can augment camera pose estimation by extending training data with rendered images. However, the images rendered by these methods are often plagued by blurring, undermining their reliability as training data for camera pose estimation. This limitation is particularly critical for Scene Coordinate Regression (SCR) methods, which aim at pixel-level 3D coordinate estimation, because rendering artifacts directly lead to estimation inaccuracies. To address this challenge, we propose a dual-criteria filtering mechanism that dynamically identifies and discards suboptimal pixels during training. The dual-criteria filter evaluates two concurrent metrics: (1) real-time SCR reprojection error, and (2) gradient threshold, across the coordinate regression domain. In addition, for visual localization problems in sparse input scenarios, it will be even more necessary to use data generated by NVS to assist the localization task. We design a coarse-to-fine PoI variant using sparse input NVS to solve this problem. Experiments across indoor and outdoor benchmarks confirm our method’s efficacy. It achieves state-of-the-art localization accuracy while maintaining computational efficiency.
@@ I Introduction
Figure 1: Left: Comparison of query and rendered images of the dataset 7Scenes and Cambridge Landmarks, revealing uneven quality within them, with query frames clear and rendered frames blurry; Right: Translation error versus training time, where ‘CoodiNet+’ denotes using rendered images as query images for CPR method CoodiNet (LENS in this case); ‘DSAC*+’ and ‘ACE+’ denote the method that combines NVS-rendered images and query images as training data for SCR method DSAC* and ACE. ‘ACE+PoI’ denotes our proposed PoI method (ACE-based); Analysis reveals that directly adding rendered data to the training set will increase training time to some extent, but performance will decrease for the SCR method. On the other hand, our PoI approach can improve the performance with an acceptable time increase.
@@
Although artifacts exist in rendered images, as shown in the left part of Figure 1, the CPR-based NRP methods achieve effectiveness mainly because they involve N-to-1 prediction (1 prediction for N pixels) and are more affected by the overall quality of the images than particular pixels. In contrast, SCR-based methods rely on accurate N-to-N prediction (N predictions for N pixels) and are particularly susceptible to the quality of rendered pixels. As shown in the right part of Figure 1, after extending the training dataset with synthetic data through NVS, the CPR method shows significant improvement, while the performance of SCR declines, accompanied by a notable increase in training time. Directly training the SCR model with raw rendered images proves less effective than CPR methods. As a result, very few studies can effectively combine NVS with the SCR framework to expand the training dataset.
@@ II Related Work
Camera Pose Regression The CPR methods, i.e., to regress the camera pose from the given image directly, are the most naive ideas in learning-based methods [20, 21, 22, 23, 24, 25, 26, 27, 12, 28]. The most straightforward method implicitly uses CNN layers or MLP to represent the image-to-pose correspondence. PoseNet [20] first proposes CPR using pre-trained GoogLeNet as the feature extractor. Then, several works focus on improving CPR through additional modules. Geomapnet [22] estimates the absolute camera poses and the relative poses between adjacent frames. AtLoc [25] uses a self-attention module to extract salient features from the image. Vlocnet++ [24] adds a semantic module to solve the dynamic scene and improve the robustness for blockings and blurs. Marepo [29] first regresses the scene-specific geometry from the input images and then estimates the camera pose using a scene-agnostic transformer. The CPR method has achieved excellent efficiency and framework simplification, but there is still room for improvement in accuracy. Scene Coordinate Regression The SCR-based method [16, 14, 30, 15, 31, 32, 13, 33] aims to estimate the camera pose in two steps: first, estimate the coordinates of the points in the 3D scene corresponding to the 2D pixels, and then estimate the camera pose based on the 2D-3D correspondence relationship. SCR was initially proposed using the random forest for RGB-D images [16]. Recently, estimating scene coordinates through RGB input has been widely studied. ForestNet[31] compares the benefits of Random Forest (RF) and Neural Networks in evaluating the scene coordinate and camera poses. ForestNet also proposes a novel method to initiate the neural network from an RF. DSAC [14], DSAC++ [30] devise a differentiable RANSAC, and thus the SCR method can be trained in an end-to-end manner. ESAC [15] uses a mixture of expert models (i.e., a gating network) to decide which domain the query belongs to, and then the complex SCR task can be split into simpler ones. DSAC* [13] extends previous works to applications using RGB-D or RGB images, with or without 3D models. This means that in the minimal case, only RGB images will be used as the input to DSAC*, just like most CPR methods. More information about the 3D structure will be utilized for most SCR methods than for CPR ones. However, approaches like DSAC* can achieve more accurate estimations even if the input is the same as the CPR method. ACE [18] and GLACE [34] abandon the time-consuming end-to-end supervision module and shuffle all pixels of the scene to improve training efficiency. ACE and GLACE utilize only RGBs without additional 3D geometry information and achieve comparable accuracy compared to former methods.
@@
Figure 2: Pipeline of our proposed methods: (a) Data Prepare: We first sample a group of synthesized camera poses $P_{novel}$ according to the query training pose $P_{query}$ using Fisher Sample. Then, we render the synthesized views $I_{novel}$ based on the sampled poses $P_{novel}$ through the novel view synthesis model. (b) Architecture of PoI module: First, a pre-trained scene-irrelevant backbone is applied to extract the features of the input query photos $I_{query}$ and the synthesized novel images $I_{novel}$. Then, the filter is applied to the features of the rendered images and extracts the features of interest. After that, we combine the query features with the retained features of the novel views and shuffle the pixel-aligned features to get the aggregation. Finally, we estimate the scene coordinates of the pixels using a scene-specific Head. The filtering algorithm is designed based on the reprojection error and the gradient of the estimated scene coordinates.
@@ III Method
In general, the pipeline of our proposed method is shown in Figure 2. For the input query images $I_{query}$ and corresponding camera poses $P_{query}$, we first sample the novel camera pose $P_{novel}$ using Fisher Sampling. Then we render novel views $I_{novel}$ through 3DGS. Finally, we use the PoI module to estimate the scene coordinates according to the input $I_{query},I_{novel}$. During the evaluation stage, we use PNP-based RANSAC to infer the camera poses from the scene coordinates.
The following part of this section is arranged as follows: We first introduce the building blocks of this paper, including NVS models, the synthetic pose sampling approach, and the baseline of the SCR regressor in Section III-A; Then we explain the details of the Fisher sample method in Section LABEL:method:fisher_sample; We elaborate on the architecture of the PoI method and introduce the details of the filtering strategy used in the PoI method in Section III-B and Section III-C; We finally explain the variant of the PoI method in extreme scenarios of sparse input in Section III-D.
@@ III-A Building Blocks
Fisher Sample Method Before PoI training, we would sample novel camera poses $P_{novel}$ and render the corresponding images $I_{novel}$ based on the query images $I_{query}$ and the corresponding camera poses $P_{query}$. Since rendered novel views are expected to contain more information about the scenes beyond the training data, we choose to use the Fisher information selection method from FisherRF [36].
@@
Backbone: To expedite convergence, we use the pseudo-depth supervision mechanism from ACE for suboptimal pixels exhibiting elevated reprojection errors. Specifically, when a pixel’s reprojection error surpasses a predefined threshold, we initially enforce its convergence toward designated pseudo-depth targets through constrained optimization. After a certain epoch, when the estimated values of these pixels fall within the valid threshold, we use conventional supervision for further optimization.
III-B Architecture of PoI
@@ III-C Filtering Strategy in PoI
$\begin{split}\mathcal{L}&=\begin{cases}\mathcal{L}_{rep}^{query}(i),&\text{if }i\in T\\ \tilde{\omega}\times\mathcal{L}_{rep}^{poi}(i),&\text{if }i\in PoI\end{cases}\\ \tilde{\omega}&=\omega_{max}-\frac{I_{iter}}{N_{iter}}(\omega_{max}-\omega_{min})\end{split}$ (2)
@@
III-D Coarse-to-Fine Strategy for Extreme Sparse Input
@@ IV Experiment
PoI(ours) 0.5/0.13 0.7/0.31 0.5/0.25 0.9/0.27 1.0/0.16 0.7/0.21 2.4/0.68 1.0/0.28
GLPoI(ours) 0.5/0.10 0.8/0.28 0.4/0.27 0.8/0.24 0.8/0.22 0.6/0.20 2.5/0.77 0.9/0.29
@@ IV-A Experiments Setup
PoI (ours) No 25min 16M 17/0.3 23/0.5 5/0.2 8/0.3 24/0.1 15/0.3
GLPoI (ours) No 25min 13M 18/0.3 15/0.4 4/0.2 8/0.3 17/0.1 12/0.3
@@ IV-D Coarse-to-Fine Experiments with Extremely Sparse Input
To further evaluate the effectiveness of our method, we conduct an additional experiment with sparse input as mentioned in Section III-D. The comparison of translation error between PoI and C2F for scene ‘fire’ is shown in Figure 7. We can see that C2F has an advantage when the input data is sparse.
@@
The numerical results are shown in Table IV, case ‘base’ denotes the sparse input circumstances with the baseline model. Case ‘coarse’ is the self-pruning step of our method using only rendered data; we still use the Fisher information sampling as the novel pose sampling method. Case ‘c2f’ denotes the fine-tuned results of our proposed method.
@@
[…]
@@
[…]